Sodium hydrogen carbonate (NaHCO₃), also known as sodium bicarbonate or "baking soda", can be used to relieve acid indigestion. Acid indigestion is the burning sensation you get in your stomach when it contains too much hydrochloric acid (HCl), which the stomach secretes to help digest food. Drinking a glass of water containing dissolved NaHCO₃ neutralizes excess HCl through this reaction:


HCl(a q)+NaHCO₃(a q) → NaCl(aq)+H₂ O(l)+CO₂(g)


The CO₂ gas produced is what makes you burp after drinking the solution.

Suppose the fluid in the stomach of a woman suffering from indigestion can be considered to be 150. mL of a 0.050 M HCl solution. What mass of NaHCO₃ would she need to ingest to neutralize this much HCl? Be sure your answer has the correct number of significant digits.
Answer #1

Moles of HCl = molarity * Volume in L = 0.05 * 0.15 = 7.5 * 10-3 moles

From the equation: moles of HCl = moles of NaHCO3 = 7.5 * 10-3 moles

Mass of NAHCO3 = moles * molar mass = 7.5 * 10-3 * (23 + 1 + 12 + 3 * 16) = 6.3 * 10-1 g